# Differential Machine Learning for 0DTE Options with Stochastic Volatility and Jumps

### 08 Mar 2026 v1

@@ 6 Conclusion
We present a differential machine learning method for 0DTE options under a stochastic-volatility jump-diffusion model. Prices are expressed in Black–Scholes form with a maturity-gated variance adjustment; training combines price-and-Greeks supervision with a PIDE-residual penalty. Across our numerical experiments, the method is substantially faster while keeping delta/vega errors low and price errors comparable. Our hedging experiment indicates that deltas produced by the network remain reliable even under stressed parameters, while gamma remains the most delicate target in the strict-short ATM regime. Moving from simulation to market data will require addressing microstructure noise and bid–ask spreads, while incorporating transaction costs and operational constraints.
@@ Appendix A Model comparison
Because 0DTE options are extremely short-dated, it is natural to ask whether simpler dynamics would suffice for the pricing and hedging checks considered in this paper. This appendix reports two model-comparison checks: (i) Black–Scholes and Merton jump-diffusion baselines versus a Bates SVJD benchmark, and (ii) an extension with contemporaneous jumps in price and variance (SVCJ) [Duffie et al.(2000)].
A.1 Black–Scholes and Merton vs a Bates benchmark
@@
We also run a one-day discrete-time $\Delta$ hedge under simulated Bates dynamics (same setup as Section 5.5) and compute hedge deltas using BS, Merton, and Bates. Table 8 reports summary statistics, both unconditionally and conditional on whether at least one jump occurs along the path. Jump realisations dominate tail outcomes and cannot be hedged away by delta hedging; differences across delta models are therefore small relative to jump-driven tail risk in this setup.
@@ A.2 SVCJ vs Bates: do volatility jumps matter for 0DTE?
To assess whether volatility jumps matter for 0DTE options, we consider an SVCJ extension (stochastic volatility with contemporaneous jumps in price and variance) in which price and variance jump simultaneously at Poisson times. The log-price jump is $Y\sim\mathcal{N}(\mu_{J},\sigma_{J}^{2})$ and the variance jump is $Z\sim\mathrm{Exp}(\text{mean}=\mu_{vJ})$, independent of $Y$ but arriving with the same intensity $\lambda$ [Duffie et al.(2000)]. We set $\mu_{vJ}=0.02$ (a positive variance-jump mean), and keep $(v_{0},\kappa,\theta,\sigma_{v},\rho,\lambda,\mu_{J},\sigma_{J})$ equal to the stressed regime above. Table 9 shows that the incremental pricing effect of adding variance jumps is small. For very short maturities, vega exposure is limited; variance jumps can therefore be weakly identified unless longer expiries are included in calibration/training. The evaluation grid here uses four same-day maturity slices, $\tau\in\{0.25,0.50,0.75,1.00\}\times(1/252)$. Accordingly, the metric “ATM” aggregates all grid points with $|x|<0.1$ across all four slices, whereas “ATM+short” additionally imposes $\tau<1/252$, i.e. it retains only the first three slices.
@@
We conduct the one-day delta-hedging under simulated SVCJ dynamics and compare hedges based on Bates vs SVCJ deltas. Table 10 shows that the P&L distributions are very similar at the reported precision.
Table 10: One-day $\Delta$-hedging P&L under simulated SVCJ dynamics: summary statistics comparing deltas from Bates vs SVCJ (negative is profit). The last column reports a nonparametric 95% bootstrap CI for CVaR1%.

### 30 Mar 2026 v2

@@ 6 Conclusion
We present a differential machine learning method for 0DTE options under a stochastic-volatility jump-diffusion model. Prices are expressed in Black–Scholes form with a maturity-gated variance adjustment; training combines joint price-and-Greek supervision with a PIDE-residual penalty. Across our numerical experiments, the method is substantially faster while keeping delta and vega errors low and price errors comparable to one-stage baselines. Our hedging experiment indicates that deltas produced by the network remain reliable even under stressed parameters, while gamma remains the most delicate target in the strict-short ATM regime. Moving from simulation to market data will require addressing microstructure noise and bid–ask spreads and incorporating transaction costs and operational constraints.
@@ Appendix A Model comparison
Because 0DTE options are extremely short-dated, it is natural to ask whether simpler dynamics would suffice for the pricing and hedging checks considered in this paper. This appendix reports two model-comparison checks: (i) Black–Scholes and Merton jump-diffusion baselines versus a Bates SVJD benchmark, and (ii) an SVCJ-type extension with contemporaneous jumps in price and variance, specified within the affine jump-diffusion framework of [Duffie et al.(2000)].
A.1 Black–Scholes and Merton versus a Bates benchmark
@@
We also run a one-day discrete-time $\Delta$ hedge under simulated Bates dynamics (using the same setup as Section 5.5) and compute hedge deltas using BS, Merton, and Bates. Table 8 reports summary statistics, both unconditionally and conditional on whether at least one jump occurs along the path. Jump realizations dominate tail outcomes and cannot be hedged away by delta hedging; differences across delta models are therefore small relative to jump-driven tail risk in this setup.
@@ A.2 SVCJ vs Bates: do volatility jumps matter for 0DTE?
To assess whether volatility jumps matter for 0DTE options, we consider an SVCJ-type extension (stochastic volatility with contemporaneous jumps in price and variance) in which price and variance jump simultaneously at Poisson times. Within the affine jump-diffusion framework of [Duffie et al.(2000)], we specify the log-price jump as $Y\sim\mathcal{N}(\mu_{J},\sigma_{J}^{2})$ and the variance jump as $Z\sim\mathrm{Exp}(\text{mean}=\mu_{vJ})$, with $Z$ independent of $Y$ but arriving with the same intensity $\lambda$. We set $\mu_{vJ}=0.02$ (a positive variance-jump mean), and keep $(v_{0},\kappa,\theta,\sigma_{v},\rho,\lambda,\mu_{J},\sigma_{J})$ equal to the stressed regime above. Table 9 shows that the incremental pricing effect of adding variance jumps is small. For very short maturities, vega exposure is limited; variance jumps can therefore be weakly identified unless longer expiries are included in calibration/training. The evaluation grid here uses four same-day maturity slices, $\tau\in\{0.25,0.50,0.75,1.00\}\times(1/252)$. Accordingly, the metric “ATM” aggregates all grid points with $|x|<0.1$ across all four slices, whereas “ATM+short” additionally imposes $\tau<1/252$, i.e. it retains only the first three slices.
@@
We conduct one-day delta hedging under simulated SVCJ dynamics and compare hedges based on Bates vs SVCJ deltas. Table 10 shows that the P&L distributions are very similar at the reported precision.
Table 10: One-day $\Delta$-hedging P&L under simulated SVCJ dynamics: summary statistics comparing deltas from Bates vs SVCJ (negative values indicate profit). The last column reports a nonparametric 95% bootstrap CI for CVaR1%.